A backyard has dimensions 3 4/5 yards by 1 1/2 yards. What is the area of the backyard in square yards (yd^2)?
Found 2 solutions by edjones, richmusick:
Answer by edjones(8007)

19/5*3/2=57/10=5.7 sq yds

Answer by richmusick(3)

Let b = 3 4/5 yds = 3.8 yds
Let h = 1 ½ yds = 1.5 yds
a = b· h
Let a = area in square yards
a = (19/5) · ( 3/2) = 57/10 = 5 7/10 yd2
or
a = (3.8) · (1.5) = 5.7 yd2